# Bayesian Concept Bottleneck Models with LLM Priors

### v1

@@ Abstract
Concept Bottleneck Models (CBMs) have been proposed as a compromise between white-box and black-box models, aiming to achieve interpretability without sacrificing accuracy. The standard training procedure for CBMs is to predefine a candidate set of human-interpretable concepts, extract their values from the training data, and identify a sparse subset as inputs to a transparent prediction model. However, such approaches are often hampered by the tradeoff between enumerating a sufficiently large set of concepts to include those that are truly relevant versus controlling the cost of obtaining concept extractions. This work investigates a novel approach that sidesteps these challenges: BC-LLM iteratively searches over a potentially infinite set of concepts within a Bayesian framework, in which Large Language Models (LLMs) serve as both a concept extraction mechanism and prior. BC-LLM is broadly applicable and multi-modal. Despite imperfections in LLMs, we prove that BC-LLM can provide rigorous statistical inference and uncertainty quantification. In experiments, it outperforms comparator methods including black-box models, converges more rapidly towards relevant concepts and away from spuriously correlated ones, and is more robust to out-of-distribution samples.
@@ 1 Introduction
Although machine learning (ML) algorithms have demonstrated remarkable predictive performance, many lack the interpretability and transparency necessary for human experts to audit their accuracy, fairness, and safety (Dwork et al.,, 2012; Rudin,, 2018). This has limited their adoption in high-stakes applications such as medicine (Thirunavukarasu et al.,, 2023) and settings where regulatory agencies require algorithms to be explainable (Meskó and Topol,, 2023).
Recent works have explored Concept Bottleneck Models (CBMs) (Koh et al.,, 2020; Kim et al.,, 2023) as a potential solution: these methods leverage black-box algorithms to extract a small number of interpretable concepts, which are subsequently processed by a fully transparent tabular model to make predictions. This allows users to understand how exactly each concept contributes to the final prediction and, if necessary, modify the extracted concepts to fix model predictions.
However, there are major limitations in the way CBMs are currently trained. They require human experts to identify a large pool of concepts a priori and annotate the concepts present in each observation. This heavy dependence on human experts is expensive and often impractical. More importantly, it is difficult to ensure this pool of concepts includes the truly relevant concepts. Consequently, resulting CBMs are often outperformed by their black-box counterparts (Yuksekgonul et al.,, 2022) and may even mislead users as to which concepts are truly relevant (Ramaswamy et al.,, 2023).
Consider the task of predicting hospital readmission from patient notes. It is impossible to prespecify all possible related concepts. Not only are there many different concepts, but any single concept can also be refined. For instance, for the concept regarding the “patient’s smoking status,” there is the more general concept regarding “substance use,” the correlated concept of “whether a patient has quit smoking,” and many more. While iterative refinement of candidate concepts may help, it would make learning a CBM even more burdensome.
LLMs have the potential to address these challenges. They are cheaper than human experts, can provide (relatively) high-quality concept annotations, and have sufficient world knowledge to hypothesize useful concepts. However, applying LLMs as a simple stand-in for a human expert does not fully address the challenges of concept discovery. First, LLM queries, though cheaper than human experts, still have a non-zero cost, which practically limits the number of concepts one can explore. Second, searching over concepts using LLMs introduces a “garden of forking paths” and may lead to the selection of spuriously correlated concepts (Gelman and Loken,, 2013). Finally, LLMs are imperfect query engines that can hallucinate, suggest bad candidate concepts, provide incorrect concept annotations, and may not even be self-consistent in their prior beliefs (Wang et al.,, 2023). As such, current methods that rely on LLMs to learn CBMs have struggled to achieve good performance, particularly in areas where the LLM’s prior knowledge is not well-aligned with the truth (Ludan et al.,, 2023; Benara et al.,, 2024).
Figure 1: Overview of Bayesian inference for learning concept bottleneck models with LLM priors (BC-LLM).
To address the aforementioned challenges, we show how wrapping LLMs within a Bayesian framework allows one to iteratively explore concepts in a way that is statistically principled and provides uncertainty quantification. An LLM can serve multiple purposes in this framework: it defines a prior over a potentially infinite set of candidate concepts, proposes concepts to explore during posterior sampling, and annotates observations with the candidate concepts. The procedure, which we refer to as BC-LLM, is designed to be computationally efficient; at each iteration, the LLM proposes a batch of candidate concepts so their respective annotations for each observation can be batched into a single LLM query. We prove that asymptotically, BC-LLM converges to the correct concepts even when the LLM prior is inconsistent or incompatible. Moreover, BC-LLM is highly generalizable and works across multiple modalities (text, images, and tabular data). Experiments across multiple datasets show that BC-LLM outperforms comparator methods and often even black-box models, converges more rapidly towards the true concepts and away from spuriously correlated ones, and is more robust to out-of-distribution (OOD) samples. Finally, when BC-LLM assisted a real-world hospital’s data science team to revise an existing tabular ML model, clinicians found the resulting CBM to be substantially more interpretable and actionable than the prior model.
@@ 2 Related work
Prior CBM methods train on observations with hand-annotated (Koh et al.,, 2020; Kim et al.,, 2023; Yuksekgonul et al.,, 2023) and, more recently, LLM-annotated concepts (Patel et al.,, 2023; McInerney et al.,, 2023; Benara et al.,, 2024) for a set of prespecified concepts by human experts. These methods are known to be highly sensitive to the selected set of concepts (Ramaswamy et al.,, 2023). More recent works have suggested using LLMs to prespecify the list of concepts instead (Oikarinen et al.,, 2023; Yang et al.,, 2023), but this requires the LLM to have accurate prior knowledge about the supervised learning task.
@@
Consequently, recent works have hypothesized that taking an iterative approach can more effectively find a set of relevant concepts (Ludan et al.,, 2023; Liu et al., 2024a, ). These methods employ a boosting-like approach, where the LLM is provided a set of misclassified examples and asked to add/revise concepts to improve performance. However, the LLM can typically only review a few observations due to limits on context length, so the resulting concept proposals are far from optimal.
@@
An alternative strategy is to conduct post-hoc distillation of a black-box model into interpretable concepts (Kim et al.,, 2018; Yeh et al.,, 2020; Zhao et al.,, 2024; Jiang et al.,, 2023), which can be used to learn an interpretable prediction model. Nevertheless, prior work has shown that the generated explanations are often not faithful to the black-box model, leading to poorer performance from the interpretable model (Doshi-Velez and Kim,, 2017).
@@
Another related line of work uses LLMs to describe conceptual differences between pairs of datasets  (Zhong et al.,, 2023, 2022; Dunlap et al.,, 2024). However, existing methods are primarily suited for simple classification tasks, as they essentially learn a CBM with a single concept rather than a set of concepts.
@@
Finally, recent works have considered combining LLMs with Bayesian techniques (Yang et al.,, 2024; Liu et al., 2024b, ; Liu et al., 2024c, ). Some have suggested that In-Context Learning (ICL) can be viewed as LLMs conducting Bayesian inference (Xie et al.,, 2021; Ye et al.,, 2024), though recent work has proven that ICL is not fully consistent with Bayesian inference (Falck et al.,, 2024). To our knowledge, this work is the first to leverage LLMs in a Bayesian framework for learning CBMs and show that rigorous statistical inference is feasible despite imperfections in LLM reasoning.
3 Method
@@
Consider a dataset $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ are model inputs and $y_{i}$ are labels. Let $\mathbf{X}\coloneq(x_{1},x_{2},\ldots,x_{n})$ and $\mathbf{y}\coloneq(y_{1},\ldots,y_{n})$. A concept $c$ defines a binary- or real-valued function $\phi_{c}$ of a model input, e.g. “Does the doctor note describe an elderly patient?” where $1$=yes and 00=no. A CBM selects a fixed number of $K$ concepts $\vec{c}=(c_{1},\ldots,c_{K})$ and fits a parametric model for $y$ using the extractions $(\phi_{c_{1}}(x),\ldots,\phi_{c_{K}}(x))$. For simplicity of presentation, our discussion will focus on logistic regression (LR)-based CBMs, although other choices are possible. In this case, the data is modeled by
$p\left(Y=1|X=x,\vec{\theta},\vec{c}\right)=\sigma\left(\sum_{k=1}^{K}\theta_{k}\phi_{c_{k}}(x)+\theta_{0}\right)$ (1)
where $\sigma$ is the sigmoid function and $\vec{\theta}=(\theta_{0},\cdots,\theta_{K})$ are the parameters. We refer to $\vec{c}$ as the “support” of the model. Throughout, we use capital and lowercase letters to differentiate between random variables and realized values, respectively. For instance, $C$ refers to a random concept whereas $c$ refers to a particular one.
@@ 3.1 Review: Bayesian variable selection
Learning CBMs shares many similarities with the more standard problem of learning sparse models in high-dimensional settings, where one must select among a finite number of features that have already been tabulated (Ishwaran et al.,, 2008; Castillo et al.,, 2015; Biswas et al.,, 2022). Here, we review the Bayesian solution for this more standard setting. In later sections, we describe extensions to the CBM setting, where there are a potentially infinite number of concepts and a cost associated with extracting each.
[…]
Inference for $p(\vec{c}|\mathbf{y},\mathbf{X})$ is typically achieved through Gibbs sampling. For each epoch, Gibbs rotates through indices $k=1,\ldots,K$, during which it replaces the $k$th concept in the current iterate $\vec{c}$ by drawing from the posterior distribution for $C_{k}$ conditional on the other concepts $\vec{c}_{-k}$. When one is unable to sample from the conditional posterior, Metropolis-within-Gibbs sampling can be used instead (Gilks et al.,, 1995). Rather than replacing the $k$th concept, it proposes a candidate concept $\check{c}_{k}$ given the other concepts per some distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ and accepts the candidate concept with probability $\alpha$. This acceptance probability $\alpha$, delineated in Algorithm 1, depends on the relative posterior probabilities of the candidate and existing concepts and is carefully designed so that the posterior distribution is stationary with respect to the sampling procedure.
@@
3.2 Bayesian variable selection with LLMs
Given the Gibbs procedures described above, there are various ways to integrate LLMs for learning CBMs. To begin, we describe two naïve approaches:
Approach 1: Propose with LLM “prior.” Let the LLM define a conditional prior over concepts, i.e. $p(C_{k}|\vec{c}_{-k})$, and propose concepts from this conditional prior, i.e. $Q(C_{k};\vec{c}_{-k})=p(C_{k}|\vec{c}_{-k})$. As concepts are generated in a data-agnostic fashion, the LLM may produce many irrelevant concepts that have low acceptance probabilities. So while Approach 1 may provide valid inference, it is inefficient and costly.
Approach 2: Propose with LLM “posterior.” Let the LLM conduct posterior inference for the $k$th concept given all the data and all but the $k$th concept through ICL and let the LLM propose from this conditional posterior, i.e. $Q(C_{k};\vec{c}_{-k})=p(C_{k}|\vec{c}_{-k},\mathbf{y},\mathbf{X})$. This simplifies Metropolis-within-Gibbs to Gibbs, which always accepts the proposed concept. As LLMs are not perfect Bayesian inference engines (Falck et al.,, 2024) and may be miscalibrated, this may lead to the inclusion of inappropriate concepts. While efficient, Approach 2 leads to biased sampling.
Instead, we propose to take the middle road, which we argue gives the best of both worlds. At each iteration, we first randomly split the data to construct subset $S$ of size $\lfloor\omega n\rfloor$ and its complement $S^{c}$, for some fraction $\omega$ (e.g. half). Then the LLM is asked to propose a candidate concept based on $\vec{c}_{-k}$ and subset $S$, which can be viewed as approximately sampling from the “partial posterior”
$Q(C_{k};\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})\approx p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X}).$ (2)
Assuming (2) holds with equality, the Metropolis acceptance probability simplifies to
$\alpha\leftarrow\min\left\{\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})},1\right\}.$
This essentially compares the likelihood of the held-out data with respect to the candidate and existing concepts, much like sample-splitting in frequentist settings. If the CBM with the candidate concept is better, it is accepted; otherwise it is accepted with probability $\alpha$. This procedure, which we refer to as Split-sample Metropolis-within-Gibbs (Algorithm 2) can learn from the data to override erroneous priors and concepts from the LLM, unlike Approach 2. In addition, it is efficient because the LLM will propose relevant concepts based on some of (though not all of) the data, unlike Approach 1.

### v2

@@ Abstract
Concept Bottleneck Models (CBMs) have been proposed as a compromise between white-box and black-box models, aiming to achieve interpretability without sacrificing accuracy. The standard training procedure for CBMs is to predefine a candidate set of human-interpretable concepts, extract their values from the training data, and identify a sparse subset as inputs to a transparent prediction model. However, such approaches are often hampered by the tradeoff between exploring a sufficiently large set of concepts versus controlling the cost of obtaining concept extractions, resulting in a large interpretability-accuracy tradeoff. This work investigates a novel approach that sidesteps these challenges: BC-LLM iteratively searches over a potentially infinite set of concepts within a Bayesian framework, in which Large Language Models (LLMs) serve as both a concept extraction mechanism and prior. Even though LLMs can be miscalibrated and hallucinate, we prove that BC-LLM can provide rigorous statistical inference and uncertainty quantification. Across image, text, and tabular datasets, BC-LLM outperforms interpretable baselines and even black-box models in certain settings, converges more rapidly towards relevant concepts, and is more robust to out-of-distribution samples.  [^1] Code for running BC-LLM and reproducing results in the paper are available at https://github.com/jjfeng/bc-llm.
@@ 1 Introduction
Although machine learning (ML) algorithms have demonstrated remarkable predictive performance, many lack the interpretability and transparency necessary for human experts to audit their accuracy, fairness, and safety [1, 2]. This has limited their adoption in high-stakes applications such as medicine [3] and settings where regulatory agencies require algorithms to be explainable [4]. Recent works have explored Concept Bottleneck Models (CBMs) [5, 6] as a potential solution: these methods leverage black-box algorithms to extract a small number of interpretable concepts, which are subsequently processed by a fully transparent tabular model to predict a target label. Thus, in principle, CBMs are much safer for use in high-stakes applications because humans can audit and, if necessary, modify the extracted concepts to fix predictions.
Nevertheless, CBMs have yet to fully realize this promise. Early CBM methods were difficult to scale because they relied on human experts to specify and annotate concepts on training data. Recent works have proposed using LLMs to suggest and even annotate concepts [7, 8, 9, 10], as LLMs are cheaper, often have sufficient world knowledge to hypothesize useful concepts, and can provide (relatively) high-quality concept annotations [9, 10]. This introduces new problems such as LLM hallucinations and inconsistencies but, more critically, does not resolve the fundamental problem that standard CBMs sacrifice a significant amount of accuracy to gain interpretability [11]. Recent proposals try to address this by encoding concepts using “soft” continuous values, such as through embedding similarity [8], rather than “hard” binary values. However, “soft” CBMs can leak information about the label that would otherwise not have been available [12, 13, 14, 15, 16]; this risk is particularly severe when concept extraction is jointly trained with the final label predictor [15]. [^2] For instance, a human would fill in 0 when asked “does this bird have a gray breast” if there is no gray breast, but a “soft” CBMs may extract 0.2 if it sees a blue breast or gray head. This provides additional hints for downstream prediction. Because the definitions of soft concepts are ambiguous, humans can no longer easily interpret, audit, or intervene on such CBMs.
The aim of this work is to minimize the accuracy gap without sacrificing the interpretability of the extracted concepts. To this end, we revisit a known pain point of CBMs: Current training procedures require a pool of candidate concepts to be identified a priori, but ensuring this pool contains all relevant concepts is difficult, particularly in domains that are yet to be scientifically well-understood and where there are an infinite number of potentially relevant concepts [17, 10]. For instance, for the task of predicting hospital readmission, there are an infinite number of factors affecting a patient’s risk. Even a single concept like “smoking status” has an infinite number of refinements, such as the correlated concept of “whether a patient has quit smoking,” the broader concept of “substance use,” and more. As shown in prior work, when the candidate pool misses important concepts, the resulting CBMs may be inaccurate and even mislead users as to which concepts are truly relevant [18].
Rather than specifying concepts upfront, we investigate iterative refinement of concepts in a CBM with the assistance of an LLM. This updates a classical modeling paradigm for the LLM era: prediction models were traditionally designed through the collaboration between domain experts and data scientists, in which they iteratively refine/engineer features for simple tabular models. LLMs can help significantly accelerate this co-design process and scale up the number of iterations, but they are also imperfect query engines that can hallucinate, suggest bad candidate concepts, incorrectly annotate concepts, and may not even be self-consistent in their prior beliefs [19]. To set up guardrails, we “wrap” the LLM within a Bayesian posterior inference procedure to explore concepts with the help of an LLM in a statistically principled manner. This approach, which we refer to as Bayesian Concept bottleneck models with LLM priors (BC-LLM), offers the following benefits:
  • BC-LLM finds relevant concepts, even in settings where there is little to no prior knowledge: Unlike prior works that soley rely on a prespecified set of candidate concepts, BC-LLM explores a potentially infinite set of concepts data-adaptively and converge towards the true ones.
  • BC-LLM is cost-effective: BC-LLM selectively annotates $O(KT)$ concepts, where $K$ is the number of concepts in the CBM and $T$ is the number of outer loops; we often use only $K\approx 5$ and $T\approx 5$. In contrast, existing methods annotate hundreds or thousands of prespecified concepts [8].
  • BC-LLM corrects LLM mistakes in a statistically rigorous manner to provide calibrated uncertainty quantification: Mistakes in the LLM’s prior are fully allowed and even expected, as they are corrected in the Metropolis accept/reject step. Moreover, BC-LLM provides uncertainty quantification of the selected concepts, which improves performance in both in-distribution and out-of-distribution (OOD) settings. We prove that BC-LLM converges to the correct concepts, asymptotically, even when the LLM prior is poor or fails to be self-consistent. The Bayesian framework also naturally protects against overfitting and overconfidence in selected concepts.
  • BC-LLM substantially improve the interpretability-accuracy tradeoff in real-world datasets: Experiments across multiple datasets and modalities (text, images, and tabular data) show that BC-LLM selects human-interpretable concepts while outperforming comparator methods, often even black-box models. Moreover, when assisting a hospital’s data science team to revise an existing tabular ML model, clinicians found BC-LLM to be substantially more interpretable and actionable.
Figure 1: BC-LLM is initialized by having the LLM hypothesize the top concepts based on keyphrases extracted from each observation (Step 0). The concepts are then iteratively refined by dropping a concept (Step 1), querying the LLM for candidate replacements (Step 2), annotating each observation with the candidate concepts using the LLM (Step 3), and determining which, if any, of the candidate concepts to accept (Step 4).
@@ 2 Related work
Prior CBM methods train on observations with hand-annotated concepts [5, 6, 20] and, more recently, concepts annotated by an LLM/VLM [21, 9, 10]. These methods are known to be highly sensitive to the selected set of concepts [18]. More recent works have suggested using LLMs to prespecify the list of candidate concepts instead [7, 8, 22, 23], but this requires the LLM to have accurate prior knowledge about the supervised learning task. We note that many of these works also make use of “soft concepts” and therefore suffer from information leakage [12].
@@
To find relevant concepts in a more data-driven manner, recent works have proposed taking an iterative approach [17, 24]. These methods employ a boosting-like approach, where the LLM is provided a set of misclassified examples and asked to add/revise concepts to improve performance. However, the LLM can typically only review a few observations due to limits on context length, so the resulting concept proposals are far from optimal.
@@
An alternative strategy is to conduct post-hoc distillation of a black-box model into interpretable concepts [25, 26, 27, 28, 29], which can be used to learn an interpretable prediction model. Nevertheless, prior work has shown that the generated explanations are often not faithful to the black-box model, leading to poorer performance from the interpretable model [30].
@@
Another related line of work uses LLMs to describe conceptual differences between pairs of datasets  [31, 32, 33]. However, existing methods are primarily suited for simple classification tasks, as they essentially learn a CBM with a single concept rather than a set of concepts.
@@
Finally, recent works have considered combining LLMs with Bayesian techniques [34, 35, 36]. Some have suggested that In-Context Learning (ICL) can be viewed as LLMs conducting Bayesian inference [37, 38], though recent work has proven that ICL is not fully consistent with Bayesian inference [39]. In contrast, this work studies how we can iteratively discover new concepts to include using LLMs using a Bayesian inference procedure, accounting for imperfections in LLM reasoning.
3 Method: BC-LLM
@@
Consider a dataset $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ are model inputs and $y_{i}$ are labels. Let $\mathbf{X}\coloneq(x_{1},x_{2},\ldots,x_{n})$ and $\mathbf{y}\coloneq(y_{1},\ldots,y_{n})$. A concept $c$ defines a binary- or real-valued function $\phi_{c}$ of a model input, e.g. “Does the doctor note describe an elderly patient?” where 11=yes and 0=no. A CBM selects a fixed number of $K$ concepts $\vec{c}=(c_{1},\ldots,c_{K})$ and fits a parametric model for $y$ using the extractions $(\phi_{c_{1}}(x),\ldots,\phi_{c_{K}}(x))$. For simplicity of presentation, our discussion will focus on logistic regression (LR)-based CBMs, although other choices are possible. In this case, the data is modeled by $p\left(Y=1|X=x,\vec{\theta},\vec{c}\right)=\sigma\left(\sum_{k=1}^{K}\theta_{k}\phi_{c_{k}}(x)+\theta_{0}\right)$ where $\sigma$ is the sigmoid function and $\vec{\theta}=(\theta_{0},\cdots,\theta_{K})$ are the parameters. We refer to $\vec{c}$ as the “support” of the model. Throughout, we use capital and lowercase letters to differentiate between random variables and realized values, respectively. For instance, $C$ refers to a random concept whereas $c$ refers to a particular one.
@@ 3.1 Review: Bayesian variable selection
Learning CBMs shares many similarities with the more standard problem of learning sparse models in high-dimensional settings, where one must select among a finite number of features that have already been tabulated [40, 41, 42]. Here, we review the Bayesian solution for this more standard setting and then discuss how it may be extended to the CBM setting.
Given priors on $\vec{c}$ and $\vec{\theta}$, the goal of Bayesian inference for sparse models in high-dimensional settings is to sample from the posterior $p(\vec{\theta},\vec{c}|\mathbf{X},\mathbf{y})$, which allows for uncertainty quantification of the true support and coefficients. Posterior samples also describe the uncertainty at a new point $x_{0}$ via the posterior predictive distribution $p(y_{0}|x_{0},\mathbf{y},\mathbf{X},\vec{c})=\iint p(y_{0}|x_{0},\vec{\theta},\vec{c})p(\vec{\theta},\vec{c}|\mathbf{y},\mathbf{X})d\vec{\theta}d\vec{c},$ which can be viewed as combining the posterior samples into an ensemble model. Factorizing the posterior per $p(\vec{\theta},\vec{c}|\mathbf{y},\mathbf{X})=p(\vec{\theta}|\vec{c},\mathbf{y},\mathbf{X})p(\vec{c}|\mathbf{y},\mathbf{X})$, the first term describes the posterior for the model parameters and the second describes the posterior over selected variables. As the former can be readily obtained by classical Bayesian inference for low-dimensional settings (e.g., posterior inference for LR coefficients), our discussion will focus on the latter.
Inference for $p(\vec{c}|\mathbf{y},\mathbf{X})$ is typically achieved through Gibbs sampling. For each outer loop, Gibbs rotates through indices $k=1,\ldots,K$, during which it replaces the $k$th concept in the current iterate $\vec{c}$ by drawing from the posterior distribution for $C_{k}$ conditional on the other concepts $\vec{c}_{-k}$. When one is unable to sample from the conditional posterior, Metropolis-within-Gibbs can be used instead [43]. Rather than immediately replacing the $k$th concept, it proposes a candidate concept $\check{c}_{k}$ given the other concepts per some distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ and accepts the candidate concept with probability $\alpha$. This acceptance probability $\alpha$, delineated in Algorithm 1, depends on the relative posterior probabilities of the candidate and existing concepts and is carefully designed so that the posterior distribution is stationary with respect to the sampling procedure.
@@
We extend Bayesian variable selection to learn CBMs, where the key difference is that the number of potential concepts is effectively infinite, which makes it difficult to formulate an appropriate prior over concepts $p(\vec{C})$ and proposal transition kernel $Q(C_{k};\vec{C}_{-k})$. Intuitively, it seems that LLMs can help perform both these tasks, but how to do so correctly and efficiently is surprisingly subtle. To motivate our choice, we first describe two naïve options:
Update 1: Propose with the LLM’s “prior.” Given the working concepts $\vec{C}_{-k}=\vec{c}_{-k}$, we can prompt the LLM to propose new concepts, without revealing any data, e.g. “Propose an additional concept for predicting readmissions, given the already-existing concepts $\vec{c}_{-k}$”.  [^3] If one believes the LLM’s prior knowledge about the outcome of interest may be misleading, we can mask this information, e.g. “Propose an additional concept for predicting some label $Y$, given the already-existing concepts $\vec{c}_{-k}$.” This can be viewed as treating the proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as sampling from the LLM’s conditional prior distribution $p(C_{k}|\vec{c}_{-k})$. Consequently, the MH acceptance ratio simplifies to $\frac{p(\mathbf{y}|(\vec{c}_{-k},\check{c}_{k}),\mathbf{X})}{p(\mathbf{y}|\vec{c},\mathbf{X})}$, which allows us to override proposed concepts that lead to poor model performance. Without access to data however, the LLM is likely to propose many irrelevant concepts from its prior, resulting in low acceptance probabilities and slow sampling. So while Update 1 may provide valid inference, it is inefficient and costly.
Update 2: Propose with the LLM’s “posterior.” Leveraging LLMs’ ability to perform ICL [37, 38], another idea is to include all the data when prompting the LLM to propose new concepts. That is, we ask the LLM itself to “conduct posterior inference” and interpret the LLM’s proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as sampling from a conditional posterior distribution $p(C_{k}|\vec{c}_{-k},\mathbf{y},\mathbf{X})$ for some implicitly defined concept prior. This approach is risky; while LLMs may be able to approximate posterior inference for small datasets, this approximation breaks down for larger sample sizes [39]. Furthermore, we can no longer override poor concepts with data as the MH acceptance ratio simplifies to 1, i.e. the proposed concept is always accepted. So while Update 2 may generate concepts more efficiently, the sampling procedure may not converge to a good or even valid posterior distribution.
These update options are at the end of two extremes, trading off between efficient sampling versus unbiased posterior inference. Our task then is to find a better balance.
3.2 The Split-sample update: Propose with the LLM’s partial posterior
Rather than relying on the LLM’s prior or posterior, a key idea in BC-LLM is to take the middle road by sampling from an LLM’s partial posterior, which we will show enjoys the best of both worlds. The proposal, which we refer to as Split-sample Metropolis update (Algorithm 2), uses the LLM to propose candidate concepts based on only a portion of the data and fixes mistakes in the LLM’s proposal using the remaining data in the accept-reject step. This addresses the drawbacks of the two naïve updating procedures in Section 3.1, as it simultaneously uses the data to generate more efficient proposals while overriding the LLM as needed.
More specifically, the split-sample update first randomly splits the data to construct subset $S$ of size $\lfloor\omega n\rfloor$ and its complement $S^{c}$, for some fraction $\omega$ (e.g. half). The LLM is prompted to generate candidate concepts by combining its prior knowledge with information from the data in $S$. We can then interpret the LLM’s proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as the (conditional) partial posterior distribution $p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$. Then in the accept-reject step, the MH acceptance probability simplifies to the partial Bayes factor $\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})}$ [44]. This compares the likelihood of the held-out data $S^{c}$ with respect to the candidate and existing concepts—much like sample-splitting in frequentist settings—and thus provides an opportunity to correct inconsistencies between the LLM’s proposal and the actual posterior distribution.